# Decentralized Proactive Model Offloading and Resource Allocation for Split and Federated Learning

### v1

@@ Abstract
In the resource-constrained IoT-edge environment, Split Federated (SplitFed) learning is implemented to enhance training efficiency. This method involves each IoT device dividing its full DNN model at a designated layer into a device-side model and a server-side model, then offloading the latter to the edge server. However, existing research overlooks four critical issues as follows: (1) the heterogeneity of IoT devices’ resource capacities and the sizes of their local data samples impact training efficiency; (2) the influence of the edge server’s computation and network resource allocation on training efficiency; (3) the data leakage risk associated with the offloaded server-side sub-model; (4) the privacy drawbacks of current centralized algorithms. Consequently, proactively identifying the optimal cut layer and server resource requirements for each IoT device to minimize training latency while adhering to data leakage risk rate constraint remains a challenging issue. To address these problems, this paper first formulates the latency and data leakage risk of training DNN models using Split Federated learning. Next, we frame the Split Federated learning problem as a mixed-integer nonlinear programming challenge. To tackle this, we propose a decentralized Proactive Model Offloading and Resource Allocation (DP-MORA) scheme, empowering each IoT device to determine its cut layer and resource requirements based on its local multidimensional training configuration, without knowledge of other devices’ configurations. Extensive experiments on two real-world datasets demonstrate that the DP-MORA scheme effectively reduces DNN model training latency, enhances training efficiency, and complies with data leakage risk constraints compared to several baseline algorithms across various experimental settings.
@@
IoT-edge Environment, Model Offloading, Resource Allocation, Data Leakage Risk, Decentralized Algorithm.
1 Introduction
@@
Various Deep Neural Networks (DNNs) have facilitated tremendous progress across a wide smart Internet of Things (IoT) applications, such as intelligent transportation [1, 2], smart healthcare [3, 4] and smart home [5]. In these smart IoT applications, large volumes of data generated by IoT devices is often private and sensitive. To utilize sensitive data to train DNN models in a safe manner, federated Learning (FL) as a privacy preserving machine learning technique, is introduced[6, 7]. FL collaborates multiple IoT devices to train a DNN model in a distributed manner while keep data locally. In FL framework, each IoT device loads a full DNN model and parallelly train its local DNN model based on local data samples, and then aggregate its local model to form a global model in an edge server. Thus, FL can obtain a DNN model without exposing sensitive data, thereby greatly satisfying data privacy requirement. However, due to the limited resource capacities of IoT devices, federated learning can suffer from two major problems: (1) IoT device with limited memory capacity could not afford to run the DNN model whose total training memory footprint exceeds the memory capacity of a single IoT device; (2) The mismatch between weak computation capacity of IoT device and prohibitive computation workload of DNN model training leads IoT-oriented FL to time-consuming and ineffective. To address these two problems, split federated (SplitFed) learning [8, 9] is introduced to split each IoT device’s full DNN model at a cut layer into a device-side model and a server-side model and offload the server-side model to the resource-adequate edge server and achieve device-edge synergy training, consequently alleviating the aforementioned two problems.
@@
Research problem about utilizing SplitFed learning to enable IoT-oriented FL to run on resource-constrained IoT devices and further improve training efficiency of IoT-oriented FL has attracted much attention in academia [8, 10, 11, 12, 13, 14]. However, existing studies mainly exist a few major problems:
@@
(1) Most studies have failed to consider that the heterogeneity of IoT devices’ resource capacities and their local data samples size on the training efficiency. If allocating the same sub-model to each IoT device, the weaker IoT device with more local data samples need to spend more time in training models, inevitably prolonging the total training latency and reducing the training efficiency of FL. For example, the authors in [8] and [14] split DNN model on each IoT device at the same cut layer and train the DNN model in parallel or sequentially.
@@
(2) Most studies have failed to consider the impact of the edge server’s computation and network resource allocation on the training efficiency. In device-edge synergy training, multiple resource-constrained IoT devices connected to an edge server for distributed DNN model training share network resource and computing resource of the edge server. An efficient resource allocation can further reduce the training latency and improve the training efficiency. Otherwise, the total training latency can be prolonged.
@@
(3) Most studies have failed to consider the data privacy leakage problem. Existing studies offload server-side sub-model to edge server from resource-constrained IoT devices, thereby accelerating the DNN model training. However, it is possible to recover raw sample data from the knowledge of the gradient parameter of offloaded server-side sub-model [15, 16]. Hence, it can incur data privacy leakage when offload server-side sub-model to server for device-edge synergy training. Therefore, it is necessary to design an appropriate model offloading strategy to trade off computation efficient of FL and its data leakage risk.
@@
(4) Most studies adopt centralized algorithms to identify the optimal cut layer for IoT devices with different resource capacities. The formulation and solution of the centralized optimization problem itself is a detriment on privacy. That is because that the centralized algorithms are generally built on the complete knowledge regarding all IoT devices’ multiple dimensional training configurations, including the computation capacity, the mini-batch size, the size of local dataset and the number of epoches, etc. Hence, it may not possible in many real-world smart applications exploiting SplitFed learning. It is more realistic for each IoT device to decide the cut layer and the required resource according to its local multiple dimensional training configuration.
@@
This paper aims to address the above problems, with a particular focus on reducing the overall training time while satisfying the data leakage risk rate constraint. We firstly present a system architecture. We then formulate the latency and data leakage risk of DNN model training adopting split federated learning. The DNN model training process consists of multiple rounds, each of which consists of Starting Phase, Intermediate Phase and End Phase three stages. We characterize the latency for these three stages and the total training latency in one training round, respectively. We next adopt a data-riven methodology to fitting the forward/backward propagation workloads, the smashed data/the smashed data’s gradient size, and the device-side model’s data size as functions of the cut layer, respectively. Based on these, we formulate the split federated learning problem as a mixed integers non-linear programming. To solve this problem, we design a decentralized proactive model offloading and resource allocation (DP-MORA) scheme which enables each IoT device to decide its own cut layer and resource requirement according to its own private information without knowing other IoT devices’ private information. Finally, extensive experimental results on real-world datasets demonstrate that compared with several baselines algorithms, the newly proposed DP-MORA scheme can reduce the total training latency while satisfying the data leakage risk rate constraint. The main contributions of this paper are summarized as follows:
@@
  • We propose a novel DP-MORA scheme. In the scheme, each IoT device can decide its own model cut layer, radio spectrum allocation and computation resource allocation according to its local multiple dimensional training configuration, without knowing the multiple dimensional training configurations of other IoT devices. The related private information of other IoT devices can be fully preserved.
@@
The rest of this paper is organized as follows. Section 2 reviews the related works. Section 3 presents the system model. Section 4 formulates joint model offloading and resource allocation problem as a mixed integers non-linear programming problem. Section 5 presents a novel DP-MORA scheme in detail. Section 6 conducts the extensive experiments and evaluates the performance of DP-MORA scheme. Finally, section 7 concludes the paper and points out the future work.
2 Related work
@@
In resource-constrained IoT-edge environment, split federated learning enables each resource-constrained IoT device to split its local full DNN model into device-side model and service-side model at the cut layer and offload the service-side model to the edge server, thereby achieving device-edge synergy training and improving training efficiency. How to identify the optimal cut layer to accelerate model training in resource-constrained IoT-edge environment has attracted much attention in academia. There are a lot of studies about exploiting split federated learning to accelerate DNN model training in a distributed manner. These studies mainly can be classified two types: efficiency-oriented split federated learning and efficiency and privacy-oriented split federated learning.
2.1 Efficiency-oriented split federated learning
@@
For efficiency-oriented device-edge synergy training, the authors [17] proposed a novel federated split learning framework to efficiently train models on distributed sequential data. Analogously, the authors in [8] designed two schemes, called SplitFedv1 and SplitFedv2, to amalgamates split learning and federated learning two approaches, and sequentially train the client/server-side models. These two schemes accelerate local training in resource constrained IoT devices by offloading partial layer of the DNN model to edge server. However, the work manually determine the same cut layer for each IoT device’s local DNN model and have failed to consider the heterogeneity of IoT devices’ resource capacities. To address the shortcoming for the identical sub-models on heterogeneous IoT devices, the authors [12] in view of the computational heterogeneity and changing network bandwidth, adopted reinforcement learning to adaptively identify the optimal cut layer for each IoT device’s local DNN model. Its main goal is to reduce the training latency. Analogously, the authors in [18] designed an approach to dynamically determine the optimal cut layer according to the state of the communication channel, and thereby improving the efficiency of model inference. The authors [13] proposed adaptive resource-aware split-learning scheme in IoT systems to trade off the training latency and energy consumption. Moreover, the authors in [19] proposed a local-loss-based split learning to optimally split the model and train the client/server-side models in parallel, aiming at to reduce the training latency. Although the aforementioned works adaptively identify the cut layers for multiple heterogeneous IoT devices to improve the training efficiency, they have failed to take the data leakage risk into account. Therefore, these solutions cannot be directly applied to data privacy requirement IoT-edge environment.
2.2 Efficiency and privacy-oriented split federated learning
@@
For an efficiency and privacy-oriented device-edge synergy training, the authors in [11] pre-train a regression model to identify the optimal cut layer and evaluate the privacy leakage risk rate by server-side model parameter quantities. To trade off the training efficiency and the data leakage risk, a federated synergy learning paradigm is proposed. However, the work only use the device-side model size to measure the privacy leakage, which lack effective demonstration in real data reconstruction attacks. To cope with this problem, the authors in [20] proposed a metric called inverse efficiency to measure privacy leakage. Based on this, they formulated model decomposition with privacy constraint in hybrid split learning and federated learning as a constraint optimization problem and transformed it into a contextual bandit problem. To address this problem, an efficient contextual bandit learning-based scheme is developed to identify the optimal cut layers for IoT devices, aiming to optimize the training latency and data privacy protection. The authors in [9] proposed a novel learning architecture that combines split learning and federated learning two approaches to improve the training efficiency and data privacy. Analogously, the authors in [21] proposed a new hybrid Federated Split Learning architecture to trade off privacy protection and the training latency. Moreover, the authors in [22] proposed to binaries the device-side model to reduce its computation workload and memory usage. To further preserve privacy, they integrated differential privacy into the binaries split learning model and trained it with additional local leak loss. Its main goal is to lightweight model and preserve privacy. Although the aforementioned works have jointly considered the cut layer selection and privacy leakage, there is still no work to consider the impact of edge server’s resource allocation on the training efficiency. Moreover, the aforementioned works adopted the centralized algorithms to identify the optimal cut layer of each individual IoT device. The formulation of the centralized optimization problem itself is a detriment on privacy.
@@
Inspired by the above motivations, we propose a metric called data leakage risk rate to measure the data leakage incurred by different selections of cut layer. Based on it, we jointly consider model offloading and resource allocation problem for split and federated learning in IoT-edge environment. To cope with this problem, we design a decentralized proactive model offloading and resource allocation scheme, aiming to minimize the training latency while satisfying the data leakage risk rate constraint.
3 System model
3.1 System architecture
@@
As illustrated in Fig. 1, we consider a device-edge synergy paradigm comprising of an edge server and $N$ heterogeneous IoT devices. These $N$ IoT devices are resource constraint and the computing capacity of each IoT device $I_{n}$ is $f_{d}^{n}$. The edge server is resource adequate and its computing capacity is $f_{s}$. By exploiting split federated learning, the edge server can cooperate with the $N$ resource-constrained IoT devices to train a global DNN model in a distributed manner without revealing their sample data to the edge server. The DNN model is composed of $L$ consecutive layers, each denoted by $l$, with $l\in\{1,...,L\}$. The layer is denoted as the fundamental component of a DNN model. The set of $N$ IoT devices can be denoted by =$\mathcal{I}=\{I_{1},...,I_{n},...,I_{N}\}$. Each IoT device $I_{n}$ owns its local dataset $\mathcal{D}_{n}=\{\textbf{z}_{n}^{i},y_{n}^{i}\}_{i=1}^{D_{n}}$, the size of which is denoted as $D_{n}=|\mathcal{D}_{n}|$. Here, $\textbf{z}_{n}^{i}\in\mathbb{R}^{1\times Q}$ and $y_{i}\in\mathbb{R}^{1\times 1}$ represent an input data sample and its corresponding label, respectively, where $Q$ denotes the dimension of the input data sample. The aggregated dataset over all IoT devices is represented by $\mathcal{D}=\bigcup_{n=1}^{N}\mathcal{D}_{n}$. A local DNN model on each IoT device $I_{n}$ is split into a device-side sub-model and a server-side sub-model at a cut layer $l_{n}=\alpha_{n}L,\alpha_{n}\in[0,1],l_{n}\in\{0,...,L\}$. The intermediate output associated with the cut layer is called smashed data. The device-side sub-model deployed on IoT device $I_{n}$ is denoted by $\mathbf{w}_{d}^{n}$. The service-side sub-model deployed on edge server is denoted by $\mathbf{w}_{s}^{n}$. The DNN model on IoT device $I_{n}$ is denoted by $\mathbf{w}_{n}=\{\mathbf{w}_{d}^{n};\mathbf{w_{s}^{n}}\}$. At each training round, the edge server cooperate these $N$ IoT devices to train their models in parallel for multiple epochs, and then aggregate their latest device-side models to the edge server to obtain the updated global DNN model. It is worth noting that the cut layer can not be the input layer for device data privacy preservation consideration. A special case is that cut layer $l_{n}=L$ means an empty server-side model. In other words, the device-edge synergy training degrades to the federated learning scheme with $N$ IoT devices. The whole training process can be divided into $\mathcal{T}=\{1,...,t,...,T\}$ rounds.
@@
3.2 The latency of DNN model training adopting split federated learning

### v2

@@ Abstract
In the resource-constrained IoT-edge computing environment, Split Federated (SplitFed) learning is implemented to enhance training efficiency. This method involves each terminal device dividing its full DNN model at a designated layer into a device-side model and a server-side model, then offloading the latter to the edge server. However, existing research overlooks four critical issues as follows: (1) the heterogeneity of end devices’ resource capacities and the sizes of their local data samples impact training efficiency; (2) the influence of the edge server’s computation and network resource allocation on training efficiency; (3) the data leakage risk associated with the offloaded server-side sub-model; (4) the privacy drawbacks of current centralized algorithms. Consequently, proactively identifying the optimal cut layer and server resource requirements for each end device to minimize training latency while adhering to data leakage risk rate constraint remains a challenging issue. To address these problems, this paper first formulates the latency and data leakage risk of training DNN models using Split Federated learning. Next, we frame the Split Federated learning problem as a mixed-integer nonlinear programming challenge. To tackle this, we propose a decentralized Proactive Model Offloading and Resource Allocation (DP-MORA) scheme, empowering each end device to determine its cut layer and resource requirements based on its local multidimensional training configuration, without knowledge of other devices’ configurations. Extensive experiments on two real-world datasets demonstrate that the DP-MORA scheme effectively reduces DNN model training latency, enhances training efficiency, and complies with data leakage risk constraints compared to several baseline algorithms across various experimental settings.
@@
IoT-Edge Computing, Model Offloading, Resource Allocation, Data Leakage Risk, Decentralized Algorithm.
I Introduction
@@
Various Deep Neural Networks (DNNs) have facilitated tremendous progress across a wide smart Internet of Things (IoT) applications, such as intelligent transportation [1, 2], smart healthcare [3, 4] and smart home [5]. In these smart IoT applications, large volumes of data generated by end devices is often private and sensitive. To utilize sensitive data to train DNN models in a safe manner, federated Learning (FL) as a privacy preserving machine learning technique, is introduced[6, 7]. FL collaborates multiple end devices to train a DNN model in a distributed manner while keeping data locally. In the FL framework, each end device loads a full DNN model and parallelly trains its local DNN model based on local data samples, and then aggregates its local model to form a global model in an edge server. Thus, FL can obtain a DNN model without exposing sensitive data, thereby greatly satisfying data privacy requirement. However, due to the limited resource capacities of end devices, federated learning can suffer from two major problems: (1) end device with limited memory capacity could not afford to run the DNN model whose total training memory footprint exceeds the memory capacity of a single end device; (2) The mismatch between weak computation capacity of the end device and prohibitive computation workload of DNN model training leads IoT-oriented FL to time-consuming and ineffective. To address these two problems, split federated (SplitFed) learning [8, 9] is introduced to split each end device’s full DNN model at a cut layer into a device-side model and a server-side model and offload the server-side model to the resource-adequate edge server and achieve device-edge synergy training, consequently alleviating the aforementioned two problems.
@@
Research problem about utilizing SplitFed learning to enable IoT-oriented FL to run on resource-constrained end devices and further improve training efficiency of IoT-oriented FL has attracted much attention in academia [8, 10, 11, 12, 13, 14]. However, existing studies mainly exist a few major problems:
@@
(1) Most studies have failed to consider the heterogeneity of end devices’ resource capacities and their local data samples size on the training efficiency. If allocating the same sub-model to each end device, the weaker end device with more local data samples need to spend more time in training models, inevitably prolonging the total training latency and reducing the training efficiency of FL. For example, the authors in [8] and [14] split the DNN model on each end device at the same cut layer and train the DNN model in parallel or sequentially.
@@
(2) Most studies have failed to consider the impact of the edge server’s computation and network resource allocation on the training efficiency. In device-edge synergy training, multiple resource-constrained end devices connected to an edge server for distributed DNN model training share network resources and computing resources of the edge server. An efficient resource allocation can further reduce the training latency and improve the training efficiency. Otherwise, the total training latency can be prolonged.
@@
(3) Most studies have failed to consider the data privacy leakage problem. Existing studies offload server-side sub-model to the edge server from resource-constrained end devices, thereby accelerating the DNN model training. However, it is possible to recover raw sample data from the knowledge of the gradient parameter of offloaded server-side sub-model [15, 16]. Hence, it can incur data privacy leakage when offloading server-side sub-model to the server for device-edge synergy training. Therefore, it is necessary to design an appropriate model offloading strategy to trade off the computation efficiency of FL and its data leakage risk.
@@
(4) Most studies adopt centralized algorithms to identify the optimal cut layer for end devices with different resource capacities. The formulation and solution of the centralized optimization problem itself is a detriment on privacy. That is because centralized algorithms are generally built on the complete knowledge regarding all end devices’ multiple dimensional training configurations, including the computation capacity, the mini-batch size, the size of the local dataset and the number of epochs, etc. Hence, it may not be possible in many real-world smart applications exploiting SplitFed learning. It is more realistic for each end device to decide the cut layer and the required resources according to its local multiple-dimensional training configurations.
@@
This paper aims to address the above problems, with a particular focus on reducing the overall training time while satisfying the data leakage risk rate constraint. We first present a system architecture. We then formulate the latency and data leakage risk of DNN model training adopting split federated learning. The DNN model training process consists of multiple rounds, each of which consists of Starting Phase, Intermediate Phase and End Phase three stages. We characterize the latency for these three stages and the total training latency in one training round, respectively. We next adopt a data-driven methodology to fit the forward/backward propagation workloads, the smashed data/the smashed data’s gradient size, and the device-side model’s data size as functions of the cut layer, respectively. Based on these, we formulate the split federated learning problem as a mixed integer non-linear programming. To solve this problem, we design a decentralized proactive model offloading and resource allocation (DP-MORA) scheme which enables each end device to decide its own cut layer and resource requirement according to its own private information without knowing other end devices’ private information. Finally, extensive experimental results on real-world datasets demonstrate that compared with several baseline algorithms, the newly proposed DP-MORA scheme can reduce the total training latency while satisfying the data leakage risk rate constraint. The main contributions of this paper are summarized as follows:
@@
  • We propose a novel DP-MORA scheme. In the scheme, each end device can decide its own model cut layer, radio spectrum allocation and computation resource allocation according to its local multiple-dimensional training configurations, without knowing the multiple-dimensional training configurations of other end devices. The related private information of other end devices can be fully preserved.
@@
The rest of this paper is organized as follows. Section 2 reviews the related works. Section 3 presents the system model. Section 4 formulates joint model offloading and resource allocation problem as a mixed integers non-linear programming problem. Section 5 presents a novel DP-MORA scheme in detail. Section 6 conducts the extensive experiments and evaluates the performance of DP-MORA scheme. Finally, section 7 concludes the paper and points out future work.
II Related work
@@
In a resource-constrained IoT-edge computing environment, split federated learning enables each resource-constrained end device to split its local full DNN model into the device-side model and server-side model at the cut layer and offload the server-side model to the edge server, thereby achieving device-edge synergy training and improving training efficiency. How to identify the optimal cut layer to accelerate model training in a resource-constrained IoT-edge computing environment has attracted much attention in academia. There are a lot of studies about exploiting split federated learning to accelerate DNN model training in a distributed manner. These studies mainly can be classified into two types: efficiency-oriented split federated learning and efficiency and privacy-oriented split federated learning.
II-A Efficiency-oriented split federated learning
@@
For efficiency-oriented device-edge synergy training, the authors [17] proposed a novel federated split learning framework to efficiently train models on distributed sequential data. Analogously, the authors in [8] designed two schemes, called SplitFedv1 and SplitFedv2, to amalgamate split learning and federated learning two approaches, and sequentially train the client/server-side models. These two schemes accelerate local training in resource-constrained end devices by offloading partial layers of the DNN model to the edge server. However, the work manually determines the same cut layer for each end device’s local DNN model and has failed to consider the heterogeneity of end devices’ resource capacities. To address the shortcomings of the identical sub-models on heterogeneous end devices, the authors [12] in view of the computational heterogeneity and changing network bandwidth, adopted reinforcement learning to adaptively identify the optimal cut layer for each end device’s local DNN model. Its main goal is to reduce the training latency. Analogously, the authors in [18] designed an approach to dynamically determine the optimal cut layer according to the state of the communication channel, and thereby improving the efficiency of model inference. The authors [13] proposed an adaptive resource-aware split-learning scheme in IoT systems to trade off the training latency and energy consumption. Moreover, the authors in [19] proposed a local-loss-based split learning to optimally split the model and train the client/server-side models in parallel, aiming at reducing the training latency. Although the aforementioned works adaptively identify the cut layers for multiple heterogeneous end devices to improve the training efficiency, they have failed to take the data leakage risk into account. Therefore, these solutions cannot be directly applied to data privacy requirement IoT-edge computing environment.
II-B Efficiency and privacy-oriented split federated learning
@@
For efficiency and privacy-oriented device-edge synergy training, the authors in [11] pre-train a regression model to identify the optimal cut layer and evaluate the privacy leakage risk rate by server-side model parameter quantities. To trade off the training efficiency and the data leakage risk, a federated synergy learning paradigm is proposed. However, the work only uses the device-side model size to measure the privacy leakage, which lacks effective demonstration in real data reconstruction attacks. To cope with this problem, the authors in [20] proposed a metric called inverse efficiency to measure privacy leakage. Based on this, they formulated model decomposition with privacy constraints in hybrid split learning and federated learning as a constraint optimization problem and transformed it into a contextual bandit problem. To address this problem, an efficient contextual bandit learning-based scheme is developed to identify the optimal cut layers for end devices, aiming to optimize the training latency and data privacy protection. The authors in [9] proposed a novel learning architecture that combines split learning and federated learning two approaches to improve the training efficiency and data privacy. Analogously, the authors in [21] proposed a new hybrid Federated Split Learning architecture to trade off privacy protection and the training latency. Moreover, the authors in [22] adopted a binarized split learning to reduce the end device’s computation workload and memory usage. To further preserve privacy, they integrated differential privacy into the binaries split learning model and trained it with additional local leak loss. Its main goal is to lightweight model and preserve privacy. Although the aforementioned works have jointly considered the cut layer selection and privacy leakage, there is still no work to consider the impact of the edge server’s resource allocation on the training efficiency. Moreover, the aforementioned works adopted the centralized algorithms to identify the optimal cut layer of each individual end device. The formulation of the centralized optimization problem itself is a detriment to privacy.
@@
Inspired by the above motivations, we propose a metric called data leakage risk rate to measure the data leakage incurred by different selections of cut layer. Based on it, we jointly consider model offloading and resource allocation problem for split and federated learning in IoT-edge computing environment. To cope with this problem, we design a decentralized proactive model offloading and resource allocation scheme, aiming to minimize the training latency while satisfying the data leakage risk rate constraints.
III System model
III-A System architecture
@@
As illustrated in Fig. 1, we consider a device-edge synergy paradigm comprising of an edge server and $N$ heterogeneous end devices. These $N$ end devices are resource constraint and the computing capacity of each end device $I_{n}$ is $f_{d}^{n}$. The edge server is resource adequate and its computing capacity is $f_{s}$. By exploiting split federated learning, the edge server can cooperate with the $N$ resource-constrained end devices to train a global DNN model in a distributed manner without revealing their sample data to the edge server. The DNN model is composed of $L$ consecutive layers, each denoted by $l$, with $l\in\{1,\cdots,L\}$. The layer is denoted as the fundamental component of a DNN model. The set of $N$ end devices can be denoted by =$\mathcal{I}=\{I_{1},\cdots,I_{n},\cdots,I_{N}\}$. Each end device $I_{n}$ owns its local dataset $\mathcal{D}_{n}=\{\textbf{z}_{n}^{i},y_{n}^{i}\}_{i=1}^{D_{n}}$, the size of which is denoted as $D_{n}=|\mathcal{D}_{n}|$. Here, $\textbf{z}_{n}^{i}\in\mathbb{R}^{1\times Q}$ and $y_{i}\in\mathbb{R}^{1\times 1}$ represent an input data sample and its corresponding label, respectively, where $Q$ denotes the dimension of the input data sample. The aggregated dataset over all end devices is represented by $\mathcal{D}=\bigcup_{n=1}^{N}\mathcal{D}_{n}$. A local DNN model on each end device $I_{n}$ is split into a device-side sub-model and a server-side sub-model at a cut layer $l_{n}=\alpha_{n}L,\alpha_{n}\in[0,1],l_{n}\in\{0,\cdots,L\}$. The intermediate output associated with the cut layer is called smashed data. The device-side sub-model deployed on end device $I_{n}$ is denoted by $\mathbf{w}_{d}^{n}$. The server-side sub-model deployed on edge server is denoted by $\mathbf{w}_{s}^{n}$. The DNN model on end device $I_{n}$ is denoted by $\mathbf{w}_{n}=\{\mathbf{w}_{d}^{n};\mathbf{w_{s}^{n}}\}$. At each training round, the edge server cooperates these $N$ end devices to train their models in parallel for multiple epochs, and then aggregate their latest device-side models to the edge server to obtain the updated global DNN model. It is worth noting that the cut layer can not be the input layer for device data privacy preservation consideration. A special case is that cut layer $l_{n}=L$ means an empty server-side model. In other words, the device-edge synergy training degrades to the federated learning scheme with $N$ end devices. The whole training process can be divided into $\mathcal{T}=\{1,\cdots,t,\cdots,T\}$ rounds. Key symbols and their descriptions used in this paper are summarized in Table I.
TABLE I: Symbols and Their Definitions
Symbol Definition
$f_{d}^{n}$ The $nth$ end device’s computing capacity
$f_{s}$ The edge server’s computing capacity
$l_{n}$ Denotes the $nth$ model’s cut layer
$D_{n}$ The $nth$ end device’s local dataset
$\Upsilon$ The local train epochs at round t
$w_{d}^{n}$ The nth device-side sub-model
$w_{s}^{n}$ The nth server-side sub-model
$r_{d,n}^{DL}$ The downlink transmission rate from the edge server to end device $I_{n}$
$\tau_{s,n}^{m,DL}$ The distribution latency of device-side sub-model $w_{d}^{n}$ for end device $I_{n}$
$\tau_{s,n}^{f,e}$ The computing latency of device-side model processing forward propagation on a mini-batch of data samples
$\tau_{d,n}^{s,UL}$ The smash data transmission latency from end device to edge server
$\tau_{s,n}^{f,e}$ The computing latency of server-side model processing a forward propagation on a single data sample
$\tau_{s,n}^{b,e}$ The computing latency of server-side model processing a backward propagation on a single data sample
$\tau_{s,n}^{g,DL}$ The smash data’s gradient transmission latency for a mini-batch data samples
$\tau_{d,n}^{b,e}$ The computing latency of device-side model processing backward propagation on a mini-batch of data samples
$\tau_{n}^{t}$ The overall training latency of end device $I_{n}$ in one training round t
$b_{d}^{n}$ The total batches of end device $I_{n}$ for an epoch
$\tau_{n,s}^{m,UL}$ The transmission latency for each end device transmits the device-side model to edge server
$\mathbf{S}_{d}^{n}(t,1)$ The smashed data gets from the end device executes its device-side model with the raw data samples
$P_{d}^{n}(l_{n})$ The data leakage risk of cut layer $l_{n}$ for each end device $I_{n}$
$\theta_{d}^{n}$ The ratio of server’s computing capacity allocated to the $nth$ end device
$\mu_{d,n}^{DL/UL}$ The ratio of communication capacity allocated to the $nth$ end device
$\alpha_{n}$ The factor for nth end device dividing the model into device-side model and server-side model
@@
III-B The latency of DNN model training adopting split federated learning